# Explicitly Modeling Subcortical Vision with a Neuro-Inspired Front-End Improves CNN Robustness

### v1

@@ C.3 Domain Shifts
Figure C4: Examples from domain shift benchmark datasets derived from ImageNet. ImageNet-Cartoon (adopted from Salvador and Oberman [salvador2022imagenetcartoon]) features object representations in cartoon style. ImageNet-Drawing (adopted from Salvador and Oberman [salvador2022imagenetcartoon]) includes images rederd as colored pencil-like hand drawings. ImageNet-Sketch (adopted from Wang et al. [wang2019learning]) consists of black-and-white sketches emphasizing contours. Stylized-ImageNet (adopted from Geirhos et al. [geirhos2022imagenettrained]) applies replaces the original textures with those of random paintings. ImageNet-R (adopted from Henrdycks et al. [hendrycks2021faces]) contains various renditions of ImageNet classes, including painting, cartoon, origami, toy, embroidery, and sculpture styles. [^2] This figure includes images adapted from multiple sources for academic illustration purposes under fair use. All images are property of their respective authors or datasets.
@@
The ImageNet-R dataset [hendrycks2021faces], consists of a curated set of 200 classes from the ImageNet validation set. This dataset includes 30,000 images featuring renditions in various artistic styles, such as paintings, sketches, and cartoons, designed to test a model’s ability to generalize beyond natural image statistics. ImageNet-R is publicly available at https://github.com/hendrycks/imagenet-r.
@@
The ImageNet-Cartoon and ImageNet-Drawing datasets [salvador2022imagenetcartoon], are two domain shift benchmarks derived from the ImageNet validation set by applying label-preserving style transformations. ImageNet-Cartoon contains images transformed into cartoon-like renditions using a GAN-based framework [9157493], while ImageNet-Drawing comprises colored pencil sketch versions of the same images created via an image processing pipeline [10.5555/2330147.2330161]. These datasets challenge models to generalize beyond natural image statistics, revealing significant accuracy drops—on average 18 and 45 percentage points, respectively—when standard ImageNet-trained models are evaluated. Both datasets are publicly available at https://zenodo.org/records/6801109 under Creative Common Attribution 4.0 International.
@@
The ImageNet-Sketch dataset [wang2019learning] is a large-scale benchmark designed to evaluate OOD generalization in image classification. It contains 50,000 black-and-white sketch-style images, with 50 images for each of the 1,000 classes in the ImageNet validation set, collected independently using keyword queries like “sketch of [class name]”. Unlike perturbation-based datasets, ImageNet-Sketch represents a significant domain shift in both texture and color, challenging models trained on natural images to rely on global structure rather than local textural cues. The dataset is publicly available at https://www.kaggle.com/datasets/wanghaohan/imagenetsketch.
@@
Stylized-ImageNet [geirhos2022imagenettrained] is created by introducing different painting styles into ImageNet images through Adaptive Instance Normalization style transfer [huang2017arbitrary]. While texture cues are replaced by those in the paintings, overall shape is preserved. Since the original dataset was introduced primarily for training purposes and models exhibited extremely low performance, we instead used a subset of Stylized-ImageNet as used in Geirhos et al. [geirhos2021partial]. This subset focuses on 16 basic categories (e.g., airplane, dog) that are supersets of 227 ImageNet classes within the WordNet hierarchy [10.1145/219717.219748]. We followed the same approach as the original article, where the probability distribution over ImageNet classes is mapped to this 16-class distribution by averaging the probabilities of corresponding fine-grained classes. This 16-class Stylized ImageNet along with the code for probability aggregation is publicly available at https://github.com/bethgelab/model-vs-human.
@@ Appendix D Additional Experiments
Given the robustness gains observed on domain shifts and image corruptions with our EVResNet50 model, we conducted a series of ablation tests to assess the contribution of individual components to this improved performance. These experiments allowed us to isolate and understand the roles of specific elements in enhancing robustness. Furthermore, to validate the adaptability of our front-end block, we integrated it into a different backbone architecture, EfficientNet-B0, and evaluated its performance across the same robustness benchmarks.
D.1 SubcorticalBlock Ablation Study
Table D1: EVResNet50 overperforms ablations on most image corruption types and on mean corruption accuracy. Clean and corrupted top-1 accuracies averaged across severities and corruptions for EVResNet50 and four ablation tests. Values indicate mean $\pm$ SD ($n=2$ seeds).
@@ D.1 SubcorticalBlock Ablation Study
EVResNet50 41.5$\pm$0.1 38.4$\pm$0.7 37.0$\pm$0.2 31.9$\pm$0.1 52.7$\pm$0.1 71.7$\pm$0.1
-- M Cells 41.4$\pm$0.2 38.6$\pm$0.1 36.7$\pm$0.2 32.2$\pm$0.1 52.3$\pm$0.2 71.7$\pm$0.0
-- Subcort. Noise 41.2$\pm$0.0 36.6$\pm$0.6 35.9$\pm$0.2 33.7$\pm$0.5 53.4$\pm$0.2 72.9$\pm$0.4
-- Light Adaptation 40.9$\pm$0.1 34.5$\pm$0.3 36.4$\pm$0.3 32.7$\pm$0.1 53.4$\pm$0.0 72.3$\pm$0.1
-- Contrast Norm. 40.6$\pm$0.3 38.1$\pm$1.0 37.4$\pm$0.1 29.3$\pm$0.0 51.5$\pm$0.1 70.4$\pm$0.2
We removed four components from the EVResNet50 model to evaluate their individual contributions: the M cell pathway, the contrast normalization layer, the light adaptation layer, and the subcortical noise. When removing subcortical noise, we ensured that the cortical noise at the end of the VOneBlock remained Poisson-distributed, consistent with the original VOneNets implementation. Despite some tradeoffs within individual corruption categories, no EVNet variants with removed components exhibited an inecrease in mean corruption accuracy (Tab. D1). As seen in VOneNets [Dapello2020.06.16.154542], stochasticity was the primary factor driving the drop in clean accuracy. Both light adaptation and contrast normalization contributed to robustness, with contrast normalization providing a slightly greater benefit. Interestingly, despite our efforts in explicitly modeling the main foveal subcortical pathways, the M cell pathway did not impact mean corruption accuracy at all, though some minor tradeoffs were observed across different corruption types.
Across all domain shift datasets, no model outperformed the full EVResNet50 (Tab. D2), although the variant without the M cell pathway achieved the same mean robustness accuracies. As with the corruption robustness results, replacing our dual noise approach with a single cortical noise injection led to a slight decrease in performance. Both normalization mechanisms contributed to OOD generalization, with light adaptation playing a more prominent role in this setting, in contrast to the corruption robustness tests where contrast normalization had a marginally greater impact.
@@
EVResNet50 38.7$\pm$0.1 57.8$\pm$0.3 33.5$\pm$0.3 38.1$\pm$0.3 23.0$\pm$0.3 41.0$\pm$0.3
-- M Cells 38.7$\pm$0.4 58.4$\pm$0.1 34.1$\pm$0.1 38.1$\pm$0.2 23.4$\pm$0.3 39.5$\pm$1.6
-- Subcort. Noise 38.2$\pm$0.3 58.0$\pm$0.2 32.2$\pm$0.4 38.1$\pm$0.0 23.9$\pm$0.4 38.8$\pm$0.6
-- Light Adaptation 38.0$\pm$0.0 56.6$\pm$0.5 33.0$\pm$0.1 38.3$\pm$0.2 23.0$\pm$0.1 39.1$\pm$0.4
-- Contrast Norm. 38.6$\pm$0.3 57.4$\pm$0.1 34.0$\pm$0.3 38.0$\pm$0.1 22.5$\pm$0.1 41.0$\pm$1.9
@@ D.2 EVNet Backend Generalization
Similarly to the results obtained with EVResNet50, the EVEfficientNet-B0 model consistently outperforms the base EfficientNet-B0 across most corruption categories, as well as in mean corruption accuracy, as shown in Table D3. However, this improvement comes with a greater relative drop in clean image accuracy compared to the ResNet50-based models. This steeper drop likely reflects architectural differences in sensitivity to input statistics. EfficientNet-B0 employs compound scaling and aggressive architecture search to optimize performance specifically for standard ImageNet inputs [tan2020efficientnet], making it more susceptible to deviations introduced by our biologically inspired preprocessing. In contrast, ResNet50, with its more generic design appears to be more adaptable to altered input distribution.
Table D3: EVEfficientNet-B0 overperforms base EfficientNet-B0 on most image corruption types and on mean corruption accuracy. Clean and corrupted top-1 accuracies averaged across severities and corruptions for EVEfficientNet-B0 and EfficientNet-B0. Values indicate mean $\pm$ SD ($n=2$ seeds).
@@
EfficientNet-B0 [tan2020efficientnet] 30.2$\pm$0.4 18.7$\pm$0.2 26.6$\pm$0.0 29.1$\pm$0.3 40.8$\pm$1.2 68.1$\pm$0.1
EVEfficientNet-B0 34.3$\pm$0.0 31.2$\pm$0.0 30.2$\pm$0.3 26.2$\pm$0.4 44.4$\pm$0.1 62.7$\pm$0.0
When evaluated on out-of-distribution datasets, EVEfficientNet-B0 surpasses the base model on all benchmarks, as reported in Table D4, which contrasts with the ResNet50 results where the base model achieved better performance on ImageNet-Sketch.
@@
EfficientNet-B0 [tan2020efficientnet] 28.9$\pm$0.2 40.2$\pm$0.4 17.1$\pm$0.5 29.9$\pm$0.2 17.3$\pm$0.5 40.0$\pm$0.7
EVEfficientNet-B0 34.2$\pm$0.1 51.1$\pm$0.3 27.8$\pm$0.1 31.8$\pm$0.0 17.5$\pm$0.0 43.3$\pm$0.1

### v2

@@ C.3 Domain Shifts
The ImageNet-R dataset [33], consists of a curated set of 200 classes from the ImageNet validation set. This dataset includes 30,000 images featuring renditions in various artistic styles, such as paintings, sketches, and cartoons, designed to test a model’s ability to generalize beyond natural image statistics. ImageNet-R is publicly available at https://github.com/hendrycks/imagenet-r.
@@
The ImageNet-Cartoon and ImageNet-Drawing datasets [35], are two domain shift benchmarks derived from the ImageNet validation set by applying label-preserving style transformations. ImageNet-Cartoon contains images transformed into cartoon-like renditions using a GAN-based framework [83], while ImageNet-Drawing comprises colored pencil sketch versions of the same images created via an image processing pipeline [84]. These datasets challenge models to generalize beyond natural image statistics, revealing significant accuracy drops—on average 18 and 45 percentage points, respectively—when standard ImageNet-trained models are evaluated. Both datasets are publicly available at https://zenodo.org/records/6801109 under Creative Common Attribution 4.0 International.
@@
The ImageNet-Sketch dataset [36] is a large-scale benchmark designed to evaluate OOD generalization in image classification. It contains 50,000 black-and-white sketch-style images, with 50 images for each of the 1,000 classes in the ImageNet validation set, collected independently using keyword queries like “sketch of [class name]”. Unlike perturbation-based datasets, ImageNet-Sketch represents a significant domain shift in both texture and color, challenging models trained on natural images to rely on global structure rather than local textural cues. The dataset is publicly available at https://www.kaggle.com/datasets/wanghaohan/imagenetsketch.
@@
Stylized-ImageNet [8] is created by introducing different painting styles into ImageNet images through Adaptive Instance Normalization style transfer [85]. While texture cues are replaced by those in the paintings, overall shape is preserved. Since the original dataset was introduced primarily for training purposes and models exhibited extremely low performance, we instead used a subset of Stylized-ImageNet as used in Geirhos et al. [7]. This subset focuses on 16 basic categories (e.g., airplane, dog) that are supersets of 227 ImageNet classes within the WordNet hierarchy [86]. We followed the same approach as the original article, where the probability distribution over ImageNet classes is mapped to this 16-class distribution by averaging the probabilities of corresponding fine-grained classes. This 16-class Stylized ImageNet along with the code for probability aggregation is publicly available at https://github.com/bethgelab/model-vs-human.
Figure C5: Examples from domain shift benchmark datasets derived from ImageNet. ImageNet-Cartoon (adopted from Salvador and Oberman [35]) features object representations in cartoon style. ImageNet-Drawing (adopted from Salvador and Oberman [35]) includes images rederd as colored pencil-like hand drawings. ImageNet-Sketch (adopted from Wang et al. [36]) consists of black-and-white sketches emphasizing contours. Stylized-ImageNet (adopted from Geirhos et al. [8]) applies replaces the original textures with those of random paintings. ImageNet-R (adopted from Henrdycks et al. [33]) contains various renditions of ImageNet classes, including painting, cartoon, origami, toy, embroidery, and sculpture styles.
@@ Appendix D Additional Experiments
D.1 EVNet Variants
Table D1: Detailed results for EVResNet50 variants, including ablations, on brain-alignment metrics. BrainScore [42, 43] V1 alignment scores and shape bias [8] for all EVResNet50 variants. ResNet50 and VOneResNet50 included for reference but not in the comparison. Values indicate mean $\pm$ SD ($n=2$ seeds).
V1 Response Properties
V1 V1 Resp. Orient. SF RF Surround Texture Resp. Resp. Shape
ResNet50 .271 .637 .822 .754 .214 .389 .792 .621 .865 18.8
±.002 ±.008 ±.027 ±.026 ±.002 ±.023 ±.028 ±.010 ±.012 ±1.2
VOneResNet50 .375 .754 .859 .969 .482 .373 .919 .792 .884 31.6
±.002 ±.006 ±.005 ±.001 ±.041 ±.003 ±.004 ±.003 ±.002 ±1.2
EVResNet50 .364 .826 .854 .950 .726 .614 .916 .781 .933 48.9
$^{\pm\text{.000}}$ $^{\pm\text{.000}}$ $^{\pm\text{.009}}$ $^{\pm\text{.000}}$ $^{\pm\text{.001}}$ $^{\pm\text{.004}}$ $^{\pm\text{.001}}$ $^{\pm\text{.000}}$ $^{\pm\text{.000}}$ $^{\pm\text{2.4}}$
- P Cells .350 .845 .854 .945 .738 .737 .910 .764 .965 77.8
$^{\pm\text{.006}}$ $^{\pm\textbf{.001}}$ $^{\pm\text{.002}}$ $^{\pm\text{.000}}$ $^{\pm\textbf{.006}}$ $^{\pm\textbf{.001}}$ $^{\pm\text{.007}}$ $^{\pm\text{.001}}$ $^{\pm\textbf{.005}}$ $^{\pm\textbf{1.7}}$
- M Cells .368 .763 .858 .950 .527 .393 .906 .781 .926 49.9
$^{\pm\text{.006}}$ $^{\pm\text{.000}}$ $^{\pm\text{.004}}$ $^{\pm\text{.000}}$ $^{\pm\text{.000}}$ $^{\pm\text{.007}}$ $^{\pm\text{.006}}$ $^{\pm\text{.004}}$ $^{\pm\text{.000}}$ $^{\pm\text{1.0}}$
- Light Adapt. .364 .826 .859 .951 .720 .630 .908 .780 .936 70.0
$^{\pm\text{.001}}$ $^{\pm\text{.001}}$ $^{\pm\text{.002}}$ $^{\pm\textbf{.000}}$ $^{\pm\text{.006}}$ $^{\pm\text{.008}}$ $^{\pm\text{.009}}$ $^{\pm\text{.005}}$ $^{\pm\text{.006}}$ $^{\pm\text{3.3}}$
- Contrast Norm. .374 .768 .868 .936 .562 .370 .921 .784 .938 48.0
$^{\pm\textbf{.007}}$ $^{\pm\text{.002}}$ $^{\pm\textbf{.002}}$ $^{\pm\text{.001}}$ $^{\pm\text{.008}}$ $^{\pm\text{.001}}$ $^{\pm\textbf{.008}}$ $^{\pm\textbf{.005}}$ $^{\pm\text{.006}}$ $^{\pm\text{2.8}}$
Table D2: Detailed results for EVResNet50 variants, including ablations, on common corruption categories. Clean and corrupted top-1 accuracies averaged across corruptions types for all EVResNet50 variants. ResNet50 and VOneResNet50 included for reference but not in the comparison. Values indicate mean $\pm$ SD ($n=2$ seeds).
@@ D.1 EVNet Variants
[…]
Table D3: Detailed results for EVResNet50 variants, including ablations, on domain-shift accuracy. Top-1 accuracies on ImageNet-{Cartoon, Drawing, R, Sketch, Stylized16} for all EVResNet50 variants. ResNet50 and VOneResNet50 included for reference but not in the comparison. Values indicate mean $\pm$ SD ($n=2$ seeds).
@@
[…]
Table D4: Detailed results for EVResNet50 variants, including ablations, on adversarial robustness. Top-1 accuracies for all EVResNet50 variants on limited adversarial set. ResNet50 and VOneResNet50 included for reference but not in the comparison. Values indicate mean $\pm$ SD ($n=2$ seeds).
$\|\delta\|_{\infty}$ $\|\delta\|_{2}$ $\|\delta\|_{1}$
Mean $\frac{1}{1020}$ $\frac{1}{255}$ 0.15 0.6 40 160
Model [%] [%] [%] [%] [%] [%] [%]
ResNet50 16.4 23.4 0.4 37.2 1.8 33.6 1.7
±0.5 ±0.8 ±0.0 ±1.0 ±0.2 ±0.7 ±0.2
VOneResNet50 50.5 62.6 30.4 66.2 42.3 64.5 37.3
±0.1 ±0.4 ±0.3 ±0.3 ±0.2 ±0.3 ±0.6
EVResNet50 53.8 62.7 38.8 65.1 48.0 64.0 44.5
$^{\pm\text{0.2}}$ $^{\pm\text{0.2}}$ $^{\pm\text{0.6}}$ $^{\pm\text{0.0}}$ $^{\pm\text{0.3}}$ $^{\pm\text{0.2}}$ $^{\pm\text{0.4}}$
- P Cells 46.7 53.4 33.4 55.6 42.4 55.2 40.3
$^{\pm\text{0.2}}$ $^{\pm\text{0.3}}$ $^{\pm\text{0.1}}$ $^{\pm\text{0.5}}$ $^{\pm\text{0.2}}$ $^{\pm\text{0.4}}$ $^{\pm\text{0.1}}$
- M Cells 54.0 62.6 40.1 64.8 48.4 64.1 44.0
$^{\pm\textbf{0.1}}$ $^{\pm\text{0.3}}$ $^{\pm\text{0.2}}$ $^{\pm\textbf{0.1}}$ $^{\pm\text{0.2}}$ $^{\pm\text{0.1}}$ $^{\pm\text{0.7}}$
- Light Adapt. 55.1 62.8 42.1 65.0 50.3 64.0 46.3
$^{\pm\text{0.5}}$ $^{\pm\textbf{0.5}}$ $^{\pm\textbf{0.4}}$ $^{\pm\text{0.7}}$ $^{\pm\text{0.9}}$ $^{\pm\text{0.2}}$ $^{\pm\textbf{0.3}}$
- Contrast Norm. 53.3 62.7 38.2 65.2 47.3 64.3 43.5
$^{\pm\text{0.1}}$ $^{\pm\text{0.2}}$ $^{\pm\text{0.4}}$ $^{\pm\textbf{0.3}}$ $^{\pm\text{0.2}}$ $^{\pm\text{0.3}}$ $^{\pm\text{0.0}}$
- Subcort. Noise 48.5 60.5 28.2 64.2 39.6 62.7 35.8
$^{\pm\text{0.1}}$ $^{\pm\text{0.1}}$ $^{\pm\text{0.8}}$ $^{\pm\text{0.1}}$ $^{\pm\text{0.3}}$ $^{\pm\text{0.1}}$ $^{\pm\text{0.3}}$
- VOneBlock 51.8 62.2 34.3 65.2 44.8 63.9 40.4
$^{\pm\text{0.5}}$ $^{\pm\text{0.3}}$ $^{\pm\text{0.6}}$ $^{\pm\textbf{0.8}}$ $^{\pm\text{0.2}}$ $^{\pm\text{0.9}}$ $^{\pm\text{0.4}}$
++ LGN-V2 Conn. 53.9 62.8 38.7 65.0 48.4 64.5 44.2
$^{\pm\text{0.2}}$ $^{\pm\textbf{0.3}}$ $^{\pm\text{0.1}}$ $^{\pm\text{0.8}}$ $^{\pm\textbf{0.2}}$ $^{\pm\textbf{0.2}}$ $^{\pm\text{0.2}}$
@@ D.2 EVNet Backend Generalization
Similarly to the results obtained with EVResNet50, the EVEfficientNet-B0 and EVCORnet-Z models consistently outperform their corresponding base model across most corruption categories, as well as in mean corruption accuracy, as shown in Table D5. However, this improvement comes with a greater relative drop in clean image accuracy compared to the ResNet50-based models. This steeper drop likely reflects architectural differences in sensitivity to input statistics. EfficientNet-B0 employs compound scaling and aggressive architecture search to optimize performance specifically for standard ImageNet inputs [5], making it more susceptible to deviations introduced by our biologically inspired preprocessing. In contrast, ResNet50, with its more generic design appears to be more adaptable to altered input distribution. Similarly, compared to ResNet50, the compact architecture of CORnet-Z exhibits a lower degree of feature redundancy which, when coupled with a mismatch between the inductive biases imposed by the front-end and those the network was designed to exploit, can limit its flexibility to adapt to the upstream processing. When evaluated on domain shift datasets, both EVEfficientNet-B0 and EVCORnet-Z surpass their base models on most benchmarks, as reported in Table D6, with the only exception being ImageNet-Sketch, mimicking the same pattern as observed with the EVResNet50. Both EVEfficientNet-B0 and EVCORnet-Z exhibit substantial improvements across all norm constraints (Table D7) and, when aggregated into the Robustness Score (Table D8), EVNets consistently surpass their respective base architectures, reinforcing the effectiveness of back-end generalization.
Table D5: EVNets outperforms base models on most image corruption types and on mean corruption accuracy, across different backend architectures. Clean and corrupted top-1 accuracies averaged across severities and corruptions for EfficientNet-B0, EVEfficientNet-B0, CORnet-Z and EVCORnet-Z. Values indicate mean $\pm$ SD ($n=2$ seeds).
@@
[…]
Table D6: EVNets outperforms base models on most OOD datasets and on mean domain shift accuracy, across different backend architectures. Top-1 accuracies on ImageNet-{Cartoon, Drawing, R, Sketch, Stylized16} for EfficientNet-B0, EVEfficientNet-B0, CORnet-Z and EVCORnet-Z. Values indicate mean $\pm$ SD ($n=2$ seeds).
@@
[…]
Table D7: EVNets outperforms base models on most adversarial perturbations and on mean adversarial robustness, across different backend architectures. Top-1 accuracies for EfficientNet-B0, EVEfficientNet-B0, CORnet-Z and EVCORnet-Z. Values indicate mean $\pm$ SD ($n=2$ seeds).
$\|\delta\|_{\infty}$ $\|\delta\|_{2}$ $\|\delta\|_{1}$
Mean $\frac{1}{1020}$ $\frac{1}{255}$ 0.15 0.6 40 160
Model [%] [%] [%] [%] [%] [%] [%]
EfficientNet-B0 20.9 35.0 2.0 43.5 5.1 36.2 3.3
$^{\pm\text{0.1}}$ $^{\pm\text{0.6}}$ $^{\pm\text{0.1}}$ $^{\pm\text{0.3}}$ $^{\pm\text{0.1}}$ $^{\pm\text{0.2}}$ $^{\pm\text{0.1}}$
EVEfficientNet-B0 45.6 53.2 31.1 55.8 40.8 55.0 37.7
$^{\pm\textbf{0.5}}$ $^{\pm\textbf{0.5}}$ $^{\pm\textbf{0.2}}$ $^{\pm\textbf{0.6}}$ $^{\pm\textbf{0.5}}$ $^{\pm\textbf{0.6}}$ $^{\pm\textbf{0.6}}$
CORnet-Z 17.6 24.4 0.7 34.8 5.4 34.7 5.5
$^{\pm\text{0.3}}$ $^{\pm\text{0.7}}$ $^{\pm\text{0.0}}$ $^{\pm\text{0.6}}$ $^{\pm\text{0.1}}$ $^{\pm\text{0.6}}$ $^{\pm\text{0.0}}$
EVCORnet-Z 31.1 36.7 19.0 38.8 26.9 39.0 26.2
$^{\pm\text{0.6}}$ $^{\pm\text{0.9}}$ $^{\pm\text{0.8}}$ $^{\pm\text{0.6}}$ $^{\pm\text{0.4}}$ $^{\pm\text{0.6}}$ $^{\pm\text{0.4}}$
Table D8: EVNets outperforms base models on Robustness Score, across different backend architectures. Robustness Score, clean and perturbed top-1 accuracies for EfficientNet-B0, EVEfficientNet-B0, CORnet-Z and EVCORnet-Z. Values indicate mean $\pm$ SD ($n=2$ seeds).
Perturbations
Robust. Score* Adversarial* Corrupt. Domain Shift Clean
Model [%] [%] [%] [%] [%]
[…]
D.3 EVNet Inference Ensembling
To evaluate whether combining the stochastic activations of EVNets across multiple forward passes leads to cumulative performance gains, we conducted an ensemble analysis varying both ensemble size and the stage at which activations are aggregated. Specifically, we compared ensembles that averaged activations at three points in the network: (1) the logit layer, (2) the final embedding stage (layer4, before the global average pooling), and (3) immediately after the VOneBlock bottleneck. We found that averaging later representations at the embedding or logit level yielded marginal but consistent improvements across clean, corruption, and domain-shift evaluations (Fig. D1). In contrast, averaging activations after the bottleneck reduced performance, with the exception of a small performance gain by the two-model ensemble when evaluated on ImageNet-C. This degradation likely arises because the bottleneck lies immediately downstream of the noise-injection stage, and averaging at this point effectively diminishes the stochastic variability that the EVNet leverages during training. We did not evaluate adversarial performance in this setting, as generating adversarial samples already requires an ensemble of forward passes, and using an additional ensemble for evaluation would compound computational costs to an impractical level.
Figure D1: Averaging EVNet logits or final embeddings yields slight performance improvements, whereas averaging bottleneck activations degrades accuracy. Accuracy under clean, corruption, and domain-shift images for EVResNet50 ensembles of varying sizes. “Logits” denotes ensembles averaged at the logit layer, “Embeddings” refers to averaging activations in layer4 (prior to global average pooling), and “Bottleneck” indicates averaging immediately after the VOneBlock bottleneck. Lines indicate mean top-1 accuracy and error bars represent SD ($n=3$ seeds)